# Multi-Excitation Projective Simulation with a Many-Body Physics-Inspired Inductive Bias

### v3

@@ Abstract
I Introduction
@@
Deep learning has become a powerful numerical tool, with various applications all over science and technology [1, 2, 3]. At the heart of this technological revolution are Artificial Neural Networks (ANN), parameterized function ansätze trained via gradient descent methods to achieve an ideal input-output behavior on data [4]. Despite the enormous success of ANNs, their complex and mostly problem-agnostic structure makes it difficult to understand their “reasoning process”, essentially turning ANNs into oracles. This, along with other pertinent issues related to reliability and trustworthiness [5, 6, 7, 8, 9], led to the development of the field known as eXplainable Artificial Intelligence (XAI) [10, 11].
@@
One promising approach to XAI realizes that many conscious human decision-making processes take the form of a chain-of-thought. The most famous machine learning approach modelling chains-of-thought is in the setting of Large Language Models (LLM)[12, 13]. The framework of Projective Simulation (PS) [14, 15] combines a model of deliberation, based on episodic memory, with reinforcement learning  [16]. It thereby extracts the essential components of chains-of-thought, and realizes that deliberation processes can be understood as a random walk of a single particle on a graph with vertices representing concepts or thoughts. Since its first proposal in [14], PS has been successfully applied to many domains [17, 18, 19, 20, 21]. In most of these applications, vertices in the graph (referred to as clips) have a more basic interpretation, e.g., as remembered percepts, or actions, or sensorimotoric memories more generally. However, the representation of chains-of-thought as simple paths in a graph is limited and cannot easily capture thoughts which are most naturally understood by taking their composite structure into account.
@@
A wide range of applications combine several concepts to arrive at new concepts. Examples include logical deductions, small arithmetic calculations, thoughts that compare the advantages and disadvantages of a potential decision, thoughts that take into account the results of early steps in the deliberation, etc. In basic PS, a single excitation/particle has to represent all the short-term information used by the agent for the current decision, not allowing it to disentangle the structure of the thoughts. Therefore, in this paper, we introduce Multi-Excitation Projective Simulation (mePS), an extension of PS to multiple particles/excitations. In this extension, the transition probabilities are allowed to depend on the full particle configuration, allowing mePS to model composite thoughts. Also here, each vertex in the graph represents an elementary concept and an excitation on a vertex expresses whether this concept is currently relevant. However, now each currently relevant concept can be represented by a separate excitation, allowing for the memory structure to be directly represented in a more disentangled fashion. Mathematically, our random walk steps now map sets of vertices to sets of vertices, naturally leading to the mathematical notion of hypergraphs [22, 23, 24].
@@
A naive implementation of mePS tends to exhibit a complexity exponential in the size of the semantic graph. The root of this exponential complexity is the fact that the size of the power set of the vertices scales exponentially with the number of vertices. Therefore, in this paper, we also present an inductive bias that reduces this complexity to a low-degree polynomial.
In machine learning, the term inductive bias [25, 26, 27] refers to restrictions or modelling assumptions imposed on the trainable models before the training starts. These restrictions can be formalizations of domain knowledge about the problem or the solution. A common example is Convolutional Neural Networks (CNN) [28, 29], which assume translation-equivariance. The restrictions can also serve the purpose of making the model easier to interpret (for example by the use of modularity [30, 31]), or making it more robust to out-of-distribution data (for example by integrating causal modeling [6, 7]).
Our inductive bias is a classical analogue of the typical structures found in many-body physics (MBP) [32, 33]. In MBP, many if not most phenomena can be understood as arising from fundamental elementary interactions of only a handful of particles. In particular, the standard model of particle physics, our most fundamental description of nature so far, only has interactions of at most four elementary particles [34, 35, 36].
@@
In this paper, we use many-body physics and its few-body interactions as inspiration for formalizing an inductive bias in classical machine learning. We prove that our inductive bias reduces the number of trainable parameters and the complexity of one random walk step from exponential to polynomial. The degree of the polynomial is given by the cutoff for how many particles are allowed to interact. Furthermore, to limit the lengths of the random walks, we introduce modifications of our inductive bias suitable for layered feed-forward hypergraphs.
@@
We numerically apply our mePS methodology and the inductive bias in three synthetic environments. The first environment is a toy model extending the Invasion Game of [14], which can be seen as a special case of contextual bandit problems [37, 38]. Here, we modify the Invasion Game to include irrelevant information, calling it the Invasion Game With Distraction. Its simplistic nature makes it a well-suited example to discuss the impact of different choices of the inductive bias. The second environment is a modification of the first with more actions and a reward that incorporates deceptive strategies used by the attacker; we call it the Deceptive Invasion Game. The final environment models the diagnosis and repair process of a broken computer, which we call the Computer Maintenance environment. In this environment, we primarily showcase the interpretability aspects of mePS agents, using the inductive bias to further illustrate the advantages of reducing agent complexity. For this purpose, we train multi-layered mePS agents. In the intermediate layer, the mePS agent hypothesizes about the causes behind observed symptoms of the malfunctioning computer before picking certain fixes that it can apply.
@@
The paper is organized as follows: First, in Section II, we describe Single-Excitation PS. In Section III, we present and define our Multi-Excitation PS agent, along with a dynamic hypergraph to model the agent’s training history in Subsection III. In Section IV, we develop the formalization of our inductive bias, including the exponential reduction in trainable parameters. However, since the motivation involves the quantum physical description of many-body systems, we will delay the presentation of this motivation towards the end in Section VI.Before, in Section V, the numerical experiments from the three learning scenarios we consider can be found in Section V. Afterwards, as already stated, we present the quantum many-body physical motivation of our classical inductive bias in Section VI.1. Based on this quantum physical foundation, we then propose approaches towards an actual quantum mePS agent in Section VI.2. Finally, in Section VII, we discuss our results and suggest some promising future directions for the mePS framework.
@@
In Appendix A, we present more details on the numerical experiments. In Appendix B, we present modifications of our inductive bias for layered hypergraphs which guarantee polynomial upper bounds on the number of steps in the random walk. Afterwards, in Appendix C, we prove that our inductive bias agents are fully compatible with the definition of mePS agents given in Section III. Then, in Appendix D, we prove the aforementioned bounds for different variants of the inductive bias.
II (Single-Excitation) Projective Simulation
@@
and initialized to 00. The glow dampening factor $\eta\in[0,1]$ is a hyperparameter, and plays a role similar to the discount factor in returns and value functions [39]. The standard PS update rule can be interpreted as a form of Hebb’s learning rule “What fires together wires together”.
III Multi-Excitation PS
While PS models chains of thought as random walks it cannot naturally represent reasoning steps that have a composite structure. For example, the decision to eat in a restaurant might depend both on the financial situation of the agent as well as their appetite. In PS, the current clip has to store all the short-term information the agent considers in the deliberation. Therefore, clips need to carry the semantics of all relevant observables, such as $c=($hungry, $\geq 100$ USD, no time to cook, good restaurant nearby)))).
@@ III Multi-Excitation PS
In addition, consider two further weight functions $g,h_{\mathrm{init}}:E\rightarrow\mathbb{R}$ for the directed hypergraph $(V,E)$, and two hyperparameters $\gamma\in[0,1]$ and $\eta\in[0,1]$. $h_{\mathrm{init}}$ gives the initialization of the h-values, $g$ gives the glow-factors or glows, $\eta$ is the glow damping factor, and $\gamma$ the forgetting factor. Before the first random walk, all glows are initialized to 00.
@@
    $h^{(n+1)}(C,C^{\prime})=$ $h^{(n)}(C,C^{\prime})-\gamma\cdot\Big{(}h^{(n)}(C,C^{\prime})-h_{\mathrm{init}}\Big{)}$
@@
If we endow $h$ with the explicit form given in Definition 6, then the entire mePS algorithm can also be viewed as a hypergraph generation tool, where hypergraphs with specific properties could be obtained after training by tailoring the agent architecture and update rule along with the learning environment. This process would produce a final hypergraph but if one also stores the generated hypergraphs at each training step, then the mePS algorithm can also generate dynamic hypergraphs, which could subsequently be analyzed using standard (hyper)graph visualization techniques [41, 42]. Much infrastructure that is currently underdeveloped in normal PS implementations, such as the single-excitation PS graph surgery rules proposed in [14], would need to first be developed before such a proposal could be fruitfully initiated. However, the inductive bias that will be described later in Section IV may provide a temporary solution to the problem of generating a hypergraph with an arbitrary size, essentially by adaptively restricting the size of the subsets appearing in the edge set $E$ during training.
@@
IV A Physics-Inspired Inductive Bias
In this section, we will present our inductive bias that will allow us to significantly reduce the computational complexity of mePS agents. Since the detailed motivation is inspired by quantum many-body physics to formulate classical analogues, we postpone a detailed explanation of the motivation to Section VI.1. Nonetheless, we emphasize that we have written said section in an accessible way which we believe to be digestible also for non-physicists.
@@ IV A Physics-Inspired Inductive Bias
    |($\Big{|}\big{(}\{c_{j_{1}},\ldots,c_{j_{i}}\}\rightarrow\{c_{k_{1}},\ldots,c_{k_{o}}\}\big{)}\in E^{(i,o)}\Big{\}}.$
@@
    Collect a list $\mathcal{H}_{\mathrm{relevant}}$ of all many-body $h$-values $h_{(i,o)}\big{(}\{c_{j_{1}},\ldots,c_{j_{i}}\},\{c_{k_{1}},\ldots,c_{k_{o}}\}\big{)}\in H^{(i,o)}$ with $(i,o)\in IO$ such that ($\big{(}\{c_{j_{1}},\ldots,c_{j_{i}}\}\rightarrow\{c_{k_{1}},\ldots,c_{k_{o}}\}\big{)}\in E^{(i,o)}$ and $\{c_{j_{1}},\ldots,c_{j_{i}}\}\subset\{c_{m_{1}},\ldots,c_{m_{x}}\}$. We refer to those as the relevant many-body h-values.
@@
    for each $h\in\mathcal{H}_{\mathrm{relevant}}$, then sample one transition ($(\{c_{j_{1}},\ldots,c_{j_{i}}\}\rightarrow\{c_{k_{1}},\ldots,c_{k_{o}}\})\in\bigcup_{(i^{\prime},o^{\prime})\in IO}E^{(i^{\prime},o^{\prime})}$.
@@
    In the original configuration $\{c_{m_{1}},\ldots c_{m_{x}}\}$, remove all excitations in $\{c_{j_{1}},\ldots,c_{j_{i}}\}$, and put excitations into $\{c_{k_{1}},\ldots,c_{k_{o}}\}$. If ($\big{(}\{c_{m_{1}},\ldots,c_{m_{x}}\}\setminus\{c_{j_{1}},\ldots,c_{j_{i}}\}\big{)}\cap\{c_{k_{1}},\ldots,c_{k_{o}}\}\neq\varnothing$, we keep those excitations and discard the second excitations for those atomic clips (see Remark 8).
@@
where $\eta\in[0,1]$ is the glow damping hyperparameter. All glows are initialized to 00.
@@
However, as we will explain in more detail in Section VI, this choice cannot naturally be linked to the behavior of any elementary particle. In fact, it is an irreversible process: the excitation that jumps on an already excited atomic clip cannot jump back and is thereby annihilated.
@@
Now, we have two hypergraphs, the ECM and the many-body hypergraph. Similarly, we have the standard $h$-values $h$ and the many-body $h$-values $h_{(i,o)}$. While conceptually related, it is not obvious that the definitions in Inductive Bias 1 are compatible with Definition 2. In Appendix C, we show that the definitions are indeed compatible during inference when using the standard probability assignment, by showing how to construct the standard $h$-values $h$ from the many-body $h$-values $h_{(i,o)}$.
@@
We are referring to this Inductive Bias with the label 2FF because in Appendix B we will introduce some small modifications of this inductive bias called 2SF and 2DP. These modifications guarantee a polynomial upper bound on the random walk path lengths. However, in the main text, we will focus on Inductive Bias 2FF because the worst-case paths are rarely encountered in practice.
Also for this inductive bias, we will see in Appendix C that it is compatible with the standard mePS agent in Definition 2 during inference with the standard probability assignment.
@@
    $h_{(2,2)}(\{c_{m},c_{n}\},\{c^{\prime}_{j},c^{\prime}_{k}\})$ such that $j,k\in\{1,2,3,4\}$, $j<k$ and $m,n\in\{1,2,3\}$, $m<n$
@@
    $h_{(2,2)}(\{c_{m},c_{n}\},\{c_{j},c^{\prime}_{k}\})$ such that $j,k\in\{1,2,3,4\}$ and $m,n\in\{1,2,3\}$, $m<n$
@@
To quantify the resource advantages provided by our inductive biases, we first consider the costs associated with an unrestricted mePS agent. For that purpose, we first make the following definition:
@@
    At each deliberation/random-walk step, there are $12^{|V|}-$ 1 relevant h-values. In particular, at each deliberation/random-walk step, one must sample from a probability distribution with $12^{|V|}-$ 1 outcomes.
@@
While Inductive Bias 2FF is naturally adapted to layered ECMs and only has an amount of trainable parameters polynomial in the number of atomic clips, it still allows for some pathological many-body hypergraphs for which there exist random walk paths that have a length exponential in the number of layers. Specifically, in Proposition 19 in Appendix D we consider many-body hypergraphs with $o>i=1$. These can start an “avalanche” of new excitations every time one tries to move one excitation forward. If one removes these excitations in the worst order (from deep to shallow), we prove that these random walk paths have an exponential length.
@@
Specifically, in Appendix B, we introduce Inductive Bias 2SF and 2DP. Inductive Bias 2SF differs from 2FF by demanding that the shallowest excitations get removed first. In Appendix D, we prove that Inductive Bias 2SF has a maximal random walk length of $\mathcal{O}(D\cdot\max_{j}{|L_{j}|})$, which is linear in both the width and depth of the ECM.
@@
Inductive Bias 2DP is the harshest one, it discards all passive excitations which did not contribute to a sampled transition. This is an agent which only keeps in mind atomic clips which are immediately relevant, but as a tradeoff its random walk lengths are upper-bounded by $\mathcal{O}(D)$, as we also prove in Appendix D.

### v4

@@ Abstract
keywords:
projective simulation, reinforcement learning, inductive bias, explainable AI, quantum-inspired
1 Introduction
@@
Deep learning has become a powerful numerical tool, with various applications all over science and technology [16, 25, 62]. At the heart of this technological revolution are Artificial Neural Networks (ANN), parameterized function ansätze trained via gradient descent methods to achieve an ideal input-output behavior on data [35]. Despite the enormous success of ANNs, their complex and mostly problem-agnostic structure makes it difficult to understand their “reasoning process”, essentially turning ANNs into oracles. This, along with other pertinent issues related to reliability and trustworthiness [44, 40, 23, 56, 69], led to the development of the field known as eXplainable Artificial Intelligence (XAI) [59, 68, 52, 60, 6, 10, 2].
@@
One promising approach to XAI realizes that many conscious human decision-making processes take the form of a chain-of-thought [66, 17]. The framework of Projective Simulation (PS) [15, 48] combines a model of deliberation, based on episodic memory, with reinforcement learning  [64]. It thereby extracts the essential components of chains-of-thought, and realizes that deliberation processes can be understood as a random walk of a single particle on a graph with vertices representing concepts or thoughts. Since its first proposal in [15], PS has been successfully applied to many domains [54, 45, 37, 50, 49]; with  [37, 49] involving the deployment of PS in real-world settings. Since PS is a special case of MEPS, all its previous applications also provide suitable application domains for MEPS. In most of these applications, vertices in the graph (referred to as clips) have a more basic interpretation, e.g., as remembered percepts, or actions, or sensorimotoric memories more generally.
@@
However, the representation of chains-of-thought as simple paths in a graph is limited and cannot easily capture thoughts which are most naturally understood by taking their composite structure into account. A wide range of applications combines several concepts to arrive at new concepts. Examples include logical deductions, small arithmetic calculations, thoughts that compare the advantages and disadvantages of a potential decision, thoughts that take into account the results of early steps in the deliberation, etc. In basic PS, a single excitation/particle has to represent all the short-term information used by the agent for the current decision, not allowing it to disentangle the structure of the thoughts. Therefore, in this paper, we introduce Multi-Excitation Projective Simulation (MEPS), an extension of PS to multiple particles/excitations. In this extension, the transition probabilities are allowed to depend on the full particle configuration, allowing MEPS to model composite thoughts. Also here, each vertex in the graph represents an elementary concept and an excitation on a vertex expresses whether this concept is currently relevant. However, now each currently relevant concept can be represented by a separate excitation, allowing for the memory structure to be directly represented in a more disentangled fashion. Mathematically, our random walk steps now map sets of vertices to sets of vertices, naturally leading to the mathematical notion of hypergraphs [14, 11, 22].
@@
A naive implementation of MEPS tends to exhibit a complexity exponential in the size of the semantic graph. The root of this exponential complexity is the fact that the size of the power set of the vertices scales exponentially with the number of vertices. Therefore, in this paper, we also present an inductive bias that reduces this complexity to a low-degree polynomial. Our inductive bias is a classical analogue of the typical structures found in many-body physics (MBP) [26, 18]. In MBP, many if not most phenomena can be understood as arising from fundamental elementary interactions of only a handful of particles. In particular, the standard model of particle physics, our most fundamental description of nature so far, only has interactions of at most four elementary particles [55, 61, 47].
@@
In this paper, we use many-body physics and its few-body interactions as inspiration for formalizing an inductive bias in classical machine learning. We prove that our inductive bias reduces the number of trainable parameters and the complexity of one random walk step from exponential to polynomial in the number of graph vertices. The degree of the polynomial is given by the cutoff for how many particles are allowed to interact. Furthermore, to limit the lengths of the random walks, we introduce modifications of our inductive bias suitable for layered feed-forward hypergraphs.
@@
We numerically apply our MEPS methodology and the inductive bias – along with a baseline comparison to agents based on standard PS, standard Q-learning, and a tabular PS-inspired multi-layer Q-learning agent – to three synthetic environments. The first environment is a toy model extending the Invasion Game of [15], which can be seen as a special case of contextual bandit problems [72, 12]. Here, we modify the Invasion Game to include irrelevant information, calling it the Invasion Game with Distraction. Its simplistic nature makes it a well-suited example to discuss the impact of different choices of the inductive bias. The second environment is a modification of the first with more actions and a reward that incorporates deceptive strategies used by the attacker; we call it the Deceptive Invasion Game. The final environment models the real-world diagnosis and repair process of a broken computer, which we call the Computer Maintenance environment. In this environment, we primarily showcase the interpretability aspects of MEPS agents, using the inductive bias to further illustrate the advantages of reducing agent complexity (over standard PS in particular). For this purpose, we train multi-layered MEPS agents. In the intermediate layer, the MEPS agent hypothesizes about the causes behind observed symptoms of the malfunctioning computer before picking certain fixes that it can apply.
@@
The paper is organized as follows: First, in Section 2, we explain how MEPS is situated within the broader scope of XAI and reinforcement learning literature. Then, in Section 3, we describe Single-Excitation PS. In Section 4, we present and define our Multi-Excitation PS agent, along with a dynamic hypergraph to model the agent’s training history in Subsection 4. In Section 5, we develop the formalization of our inductive bias, including the exponential reduction in trainable parameters. However, since the motivation involves the quantum physical description of many-body systems, we will delay the presentation of this motivation towards the end in Section 7.Before, in Section 6, the numerical experiments from the three learning scenarios we consider can be found in Section 6. Afterwards, as already stated, we present the quantum many-body physical motivation of our classical inductive bias in Section 7.1. Based on this quantum physical foundation, we then propose approaches towards an actual quantum MEPS agent in Section 7.2. Finally, in Section 8, we discuss our results and suggest some promising future directions for the MEPS framework.
@@
In  A, we present more details on the numerical experiments. In  B, we present modifications of our inductive bias for layered hypergraphs which guarantee polynomial upper bounds on the number of steps in the random walk. Afterwards, in C, we prove that our inductive bias agents are fully compatible with the definition of MEPS agents given in Section 4. Then, in  D, we prove the aforementioned bounds for different variants of the inductive bias.
2 Related Works
The field of XAI [59, 68, 52, 60, 6, 10] is dedicated to making AI systems more human-understandable. One can broadly distinguish three different approaches. Concerning the success of ANNs, model-agnostic explanations seek to make the decisions of artificial neural networks more understandable without affecting the architecture or training of the model. A large focus of the research community is on feature importance methods, such as Shapley values [58]. These measure the strength with which a part of the input or of the model contributes to the decision. Another commonly mentioned method is given by Local Interpretable Model-agnostic Explanations (LIME) [38]. These methods consider an environment of a data point, and within this environment they approximate the model with a much simpler model, e.g. a linear approximation. Another important class of explanation methods within the family of model-agnostic interpretations is given by so-called counterfactual explanations. A typical use-case is the formulation of “What if something were different?” questions to see how hypothetical changes would influence the model’s decisions [51].
A second approach to XAI is specific to Large Language Models (LLM). These models can communicate with humans in natural language, providing an inherent advantage concerning their understandability. Crucially, the term chain of thought was popularized within this setting [66, 17]. The basic idea is to explicitly prompt a LLM to provide a step-by-step reasoning process to answer a question formulated with natural language. This provides a human-understandable justification for an answer. Furthermore, it turned out that the accuracy of the final answer itself also increased. MEPS takes a very different approach to chains of thought, not relying on natural language and the associated expensive costs, in particular regarding big data. Instead, the idea is to design a problem-specific hypergraph in such a way that a random walk path helps the user understand how to solve a problem. The excitations then can be interpreted as the currently relevant hypergraph semantics, while a random walk step is a formalization of the thought transitioning to the next relevant semantics embedded into the hypergraph design.
The last broad class of approaches to XAI is given by mechanistic interpretability [41]. This class seeks to obtain a detailed understanding of the individual steps of decision making processes, e.g. by discovering analogous sub-circuits [20]. The extreme form are models with inherent interpretability in which the architecture is designed right from the start to allow humans to understand its decision making process. Commonly mentioned examples are linear models [38] and decision trees [21]. Causal models [40, 23] intrinsically model environments and agents using cause-effect relations of observables, and promise to overcome harmful spurious correlations in data. The field of NeuroSymbolic AI (NeSy) [70, 34] works towards combining the rigorous explanation power of mathematical logic with the efficiency of ANN methods such as gradient descent. At last, also PS and our MEPS belong into the category of intrinsically interpretable models.
Beyond the endeavour to use explainable AI to increase human understanding of the application domain, XAI is commonly mentioned together with the goal of trustworthiness [24, 2]. As AI becomes more integrated into human activities involving decision-making that impacts the lives of other people, it is important that we can trust these systems to make ethically responsible choices we would normally entrust to humans – even more so since many of these systems are opaque at their foundation. Constructing transparent agent architectures should be the first step towards helping human experts develop machine learning models which are inherently more trustworthy. It is the purpose of this work to introduce such an architecture in the form of MEPS, leaving the philosophical considerations associated with the notion of trustworthiness to other experts for the time being [25, 2, 24, 32, 6, 10].
Both PS and the MEPS methodology introduced in this work are qualitatively very different from ANNs. The discrete nature of the (hyper)graph, the random walks and the update rule makes MEPS a tabular method. A standard example for tabular methods are Q-learning agents [64]. Its tabular nature means that MEPS inherits the associated scalability problems also afflicting Q-learning. However, Q-learning only provides a look-up table for observation-action pairs, but does not provide a refined decision-making process involving intermediate steps. In particular, tabular Q-learning only provides interpretations in the form of a lookup table for reflex agents. The fact that MEPS is not a reflex agent is reflected in the fact that MEPS agents can be modelled to have several tabular layers, instead of just one, and that a decision-making process takes several random walk steps. In particular, such a refined layered structure gives several avenues for counterfactual interpretations specific to MEPS. The sampling in intermediate layers means that the agent decides that certain atomic clips (and their associated semantics) are relevant for the current deliberation, when it could have picked other atomic clips. One can even counterfactually alter the decision-making process by manually adding or removing excitations during a deliberation. Note that Q-learning can also be extended beyond the reflex agent paradigm by using PS as an inspiration to define a notion of multi-layered Q-learning agents (see Subsection 6.3 for details).
Besides explainability, another key topic in our paper is given by inductive biases. In machine learning, the term inductive bias [36, 46, 57] refers to restrictions or modelling assumptions imposed on the trainable models before the training starts. These restrictions can be formalizations of domain knowledge about the problem or the solution. A common example is Convolutional Neural Networks (CNN) [1, 19], which assume translation-equivariance. The restrictions can also serve the purpose of making the model easier to interpret (for example by the use of modularity [63, 4]), or making it more robust to out-of-distribution data (for example by integrating causal modeling [40, 23]). However, our inductive biases are qualitatively very different, because they are tailored to MEPS rather than ANNs.
3 (Single-Excitation) Projective Simulation
@@ 2 Related Works
and initialized to 0. The glow dampening factor $\eta\in[0,1]$ is a hyperparameter, and plays a role similar to the discount factor in returns and value functions [13]. The standard PS update rule can be interpreted as a form of Hebb’s learning rule “What fires together wires together”.
4 Multi-Excitation PS
While PS models chains of thought as random walks it cannot naturally represent reasoning steps that have a composite structure. For example, the decision to eat in a restaurant might depend both on the financial situation of the agent as well as their appetite. In PS, the current clip has to store all the short-term information the agent considers in the deliberation. Therefore, clips need to carry the semantics of all relevant observables, such as c=(c=(hungry, $\geq 100$ USD, no time to cook, good restaurant nearby)).
@@ 4 Multi-Excitation PS
In addition, consider two further weight functions $g,h_{\mathrm{init}}:E\rightarrow\mathbb{R}$ for the directed hypergraph $(V,E)$, and two hyperparameters $\gamma\in[0,1]$ and $\eta\in[0,1]$. $h_{\mathrm{init}}$ gives the initialization of the h-values, $g$ gives the glow-factors or glows, $\eta$ is the glow damping factor, and $\gamma$ the forgetting factor. Before the first random walk, all glows are initialized to 0.
@@
    $h^{(n+1)}(C,C^{\prime})=\$$h^{(n)}(C,C^{\prime})-\gamma\cdot\Big(h^{(n)}(C,C^{\prime})-h_{\mathrm{init}}\Big)$
@@
If we endow $h$ with the explicit form given in Definition 6, then the entire MEPS algorithm can also be viewed as a hypergraph generation tool, where hypergraphs with specific properties could be obtained after training by tailoring the agent architecture and update rule along with the learning environment. This process would produce a final hypergraph but if one also stores the generated hypergraphs at each training step, then the MEPS algorithm can also generate dynamic hypergraphs, which could subsequently be analyzed using standard (hyper)graph visualization techniques [8, 29]. Much infrastructure that is currently underdeveloped in normal PS implementations, such as the single-excitation PS graph surgery rules proposed in [15], would need to first be developed before such a proposal could be fruitfully initiated. However, the inductive bias that will be described later in Section 5 may provide a temporary solution to the problem of generating a hypergraph with an arbitrary size, essentially by adaptively restricting the size of the subsets appearing in the edge set $E$ during training.
@@
5 A Physics-Inspired Inductive Bias
In this section, we will present our inductive bias that will allow us to significantly reduce the computational complexity of MEPS agents. Since the detailed motivation is inspired by quantum many-body physics to formulate classical analogues, we postpone a detailed explanation of the motivation to Section 7.1. Nonetheless, we emphasize that we have written said section in an accessible way which we believe to be digestible also for non-physicists.
@@ 5 A Physics-Inspired Inductive Bias
    $\Big|\big(\{c_{j_{1}},\ldots,c_{j_{i}}\}\rightarrow\{c_{k_{1}},\ldots,c_{k_{o}}\}\big)\in E^{(i,o)}\Big\}.$
@@
    Collect a list $\mathcal{H}_{\mathrm{relevant}}$ of all many-body $h$-values $h_{(i,o)}\big(\{c_{j_{1}},\ldots,c_{j_{i}}\},\{c_{k_{1}},\ldots,c_{k_{o}}\}\big)\in H^{(i,o)}$ with $(i,o)\in IO$ such that $\big(\{c_{j_{1}},\ldots,c_{j_{i}}\}\rightarrow\{c_{k_{1}},\ldots,c_{k_{o}}\}\big)\in E^{(i,o)}$ and $\{c_{j_{1}},\ldots,c_{j_{i}}\}\subset\{c_{m_{1}},\ldots,c_{m_{x}}\}$. We refer to those as the relevant many-body h-values.
@@
    for each $h\in\mathcal{H}_{\mathrm{relevant}}$, then sample one transition $(\{c_{j_{1}},\ldots,c_{j_{i}}\}\rightarrow\{c_{k_{1}},\ldots,c_{k_{o}}\})\in\bigcup_{(i^{\prime},o^{\prime})\in IO}E^{(i^{\prime},o^{\prime})}$.
@@
    In the original configuration $\{c_{m_{1}},\ldots c_{m_{x}}\}$, remove all excitations in $\{c_{j_{1}},\ldots,c_{j_{i}}\}$, and put excitations into $\{c_{k_{1}},\ldots,c_{k_{o}}\}$. If $\big(\{c_{m_{1}},\ldots,c_{m_{x}}\}\setminus\{c_{j_{1}},\ldots,c_{j_{i}}\}\big)\cap\{c_{k_{1}},\ldots,c_{k_{o}}\}\neq\varnothing$, we keep those excitations and discard the second excitations for those atomic clips (see Remark 8).
@@
where $\eta\in[0,1]$ is the glow damping hyperparameter. All glows are initialized to 0.
@@
However, as we will explain in more detail in Section 7, this choice cannot naturally be linked to the behavior of any elementary particle. In fact, it is an irreversible process: the excitation that jumps on an already excited atomic clip cannot jump back and is thereby annihilated.
@@
Now, we have two hypergraphs, the ECM and the many-body hypergraph. Similarly, we have the standard $h$-values $h$ and the many-body $h$-values $h_{(i,o)}$. While conceptually related, it is not obvious that the definitions in Inductive Bias 1 are compatible with Definition 2. In C, we show that the definitions are indeed compatible during inference when using the standard probability assignment, by showing how to construct the standard $h$-values $h$ from the many-body $h$-values $h_{(i,o)}$.
@@
We are referring to this Inductive Bias with the label 2FF because in  B we will introduce some small modifications of this inductive bias called 2SF and 2DP. These modifications guarantee a polynomial upper bound on the random walk path lengths. However, in the main text, we will focus on Inductive Bias 2FF because the worst-case paths are rarely encountered in practice.
Also for this inductive bias, we will see in  C that it is compatible with the standard MEPS agent in Definition 2 during inference with the standard probability assignment.
@@
    $h_{(2,2)}(\{c_{m},c_{n}\},\{c^{\prime}_{j},c^{\prime}_{k}\})$ such that $\in\{1,2,3,4\}, j<kj<k$ and $\in\{1,2,3\}, m<nm<n$
@@
    $h_{(2,2)}(\{c_{m},c_{n}\},\{c_{j},c^{\prime}_{k}\})$ such that $j,k\in\{1,2,3,4\}$ and $\in\{1,2,3\}, m<nm<n$
@@
To quantify the resource advantages provided by our inductive biases, we first consider the costs associated with an unrestricted MEPS agent, which can be thought of as a standard PS agent by way of mapping the graph-based ECM to the hypergraph ECM defined in this work. For that purpose, we first make the following definition:
@@
    At each deliberation/random-walk step, there are $2^{|V|}-1$ relevant h-values. In particular, at each deliberation/random-walk step, one must sample from a probability distribution with $2^{|V|}-1$ outcomes.
@@
While Inductive Bias 2FF is naturally adapted to layered ECMs and only has an amount of trainable parameters polynomial in the number of atomic clips, it still allows for some pathological many-body hypergraphs for which there exist random walk paths that have a length exponential in the number of layers. Specifically, in Proposition 19 in  D we consider many-body hypergraphs with $o>i=1$. These can start an “avalanche” of new excitations every time one tries to move one excitation forward. If one removes these excitations in the worst order (from deep to shallow), we prove that these random walk paths have an exponential length.
@@
Specifically, in  B, we introduce Inductive Bias 2SF and 2DP. Inductive Bias 2SF differs from 2FF by demanding that the shallowest excitations get removed first. In  D, we prove that Inductive Bias 2SF has a maximal random walk length of $\mathcal{O}(D\cdot\max_{j}{|L_{j}|})$, which is linear in both the width and depth of the ECM.
@@
Inductive Bias 2DP is the harshest one, it discards all passive excitations which did not contribute to a sampled transition. This is an agent which only keeps in mind atomic clips which are immediately relevant, but as a tradeoff its random walk lengths are upper-bounded by $\mathcal{O}(D)$, as we also prove in  D.